Summary for DAY shift, Nov. 12, 1999

Angela Bellavance (bellavan@rice.edu)
Fri, 12 Nov 1999 15:29:44 -0600 (CST)

Shift Summary for day shift, 12/Nov/99

On Shift: J. Graham, A. Bellavance

Hours of beam delivered at all intensities: 2.4
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 2.4 (0.7 hrs usable)

Accounting of lost time: Recovering from TeV down time; many TeV problems

========================================================================
Runs ended this shift:
Run 14950 - ? evts - BEAM_MUON - no good. targeting angle off
Run 14951 - on-going - BEAM_MUON

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none
========================================================================
Overview of shift:
Waiting for beam; attempt to do a muon run with spotty beam. Lots of
TeV problems.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
7:00 Still same message on channel 13: "Trouble with a
Booster Correction Element power supply. Experts en
route."
7:29 Ch. 13: "Beam in TeV, establishing beam to switchyard
dump. Problems with MUSEP continue. Estimate beam
to users, 30 minutes."
8:09 Ch. 13: "Problems with MUSEP continue. Preparing for
switchyard access. Estimate beam to users: 60-90
minutes."
9:58 Ch. 13: "Preparing to send beam to users."
10:27 Ch. 13: "S:Q101 is holding off beam. Investigating."
10:59 Ch. 13: "Estimate time of beam to users: 20 mins."
11:28 Ch. 13: "Estimate time of beam to users: 20 mins."
14950 1 11:34 We get one crappy spill of beam!
" 11:35 MCR calls to warn us beam is coming, but it has
" disappeared.
" 11:50 Ch. 13: "Experiencing TLG problems."
" 12:06 Beam returns but targeting is way off. MCR notified,
" but they are still working on TeV problems.
" 12:08 Ch. 13: "Intermittent beam to users."
14950 13:03 Beam is so crappy and sporatic we decide to stop
the muon run and start a new one. While stopping the
run, beam goes away.
13:11 Ch. 13: "Q101 problems. Experts investigating."
14951 1 13:38 Beam returns to trick us into starting a muon run,
" then disappears again.
" ~14:00 Beam returns, pretty good targeting, but
" intermittent and low intensity.
" 14:19 Ch. 13: "Beam to users; may be intermittent due to
" problems with Q101.."
14951 40 Beam more stable now but not collecting very many
" events.
" 52 Beam disappears.
" 14:46 Ch. 13: "Q101 problems. Experts investigating."
" 15:00 Ch. 13: "Estimate 30 minutes for beam."
" 15:30 Hand over to evening crew. No beam yet.

========================================================================
Follow up Notes on Previous Shift

Muon run in progress.
Magnet is ready to turn on to + (but is still off).
Ped. run not done yet.

========================================================================
Items to Follow up on Next Shift

Ramp up analysis magnet.
Do a pedistal run and install new constants.
Take data.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke